OpenStudy (anonymous):

Hi...given this 4th degree trinomial: x4 + x2 +1, factor complete into product of two second degree trinomials and check(to obtain x4+x2+1). Please help, thank you.

OpenStudy (anonymous):

x4 + x2 +1 =x4 + 2x2 +1 -x2 =(x^2+1)^2 - x^2 =(x^2+1 +x)(x^2+1-x)

OpenStudy (anonymous):

I've noticed you've expressed the exponents a bit different. Does this affect the answer? Just the same....thank you so much.

OpenStudy (anonymous):

just the same no difference welcome